Information on pending and completed system maintenance issues
7/7/2023 -- Significant damage to the observatory due to lightening. Some of the damage has been repaired and some is pending repair. The system is offline at the moment.
2/19/2023 -- Outage at AnyDesk corporate -- cannot connect to MPO computers - resolved
2/21/2023 -- Ongoing imaging efforts at MPO and at other AAVSO observatories.
10/14/2022 -- Investigating the possibility of supporting a UT program re young star planet formation.
8/3/2022 -- Control circuit is functional. Returning the 10-slot filter wheel to service will require a trip to the site, which is not yet scheduled.
6/18/2022 -- Additional failure of the 10 slot filter wheel on observatory -- the control circuit board inside the wheel housing failed and is pending repair (we hope). We have installed a 7 slot filter wheel (on loan from Arne Henden) and reconfigured the filter set. (See the 'Telescope Specs' page for details). Installed a KVM (keyboard, video, mouse) switch to accommodate 2 computers. The system successfully ran last night.
5/16/2022 -- Malfunction of filter wheel had observatory off line for a few weeks. It has been repaired and the observatory is back on line and operating normally.
1/26/2022 -- Installed repaired camera on telescope. Replaced the camera rotator with a spare to return the system to service. Greased the dome wheels. Replaced various observatory red LEDs. Changed three of the filters in the filter wheel (see Specifications page).
11/19/2021 -- Problem with the Dec encoder was resolved by replacing the malfunctioning one with a spare. Meanwhile, a problem was found with the RA encoder and resolved. Camera is in the process of being shipped for repair. Spare camera is installed on the telescope. One of the two computers is working properly with a new video card and monitor. The other computer is a small form factor and no suitable video card has yet been installed.
A new capability at the observatory is in place -- a set of cameras that monitors and reports meteor passes. This is part of a system of meteor monitors deployed by amateur astronomers.
9/21-22/2021 -- Telescope work -- Install 'backup' imaging camera. \ More work to be done remotely to establish the backup camera service, Monitor issues (lightning strike) -- Both the obs computer and the wx computer are functional, but both of these computers have no display on the monitor. The two monitors in the observatory are damaged (these are the older Dell monitors). We temporarily hooked up one of the monitors (known to be working) from the house computer and this didn't work on either the obs or the wx computer. It seems that the video circuitry in the computers is damaged. The computers seem to be otherwise ok. Both computers have a PCI-E bus, so we'll need to purchase two new (not-high-end) video cards and new HDMI monitors. Ground Fault Issues -- The UPS units report 'ground fault problems', which would indicate that there's some voltage between ground and power common. These are very short term events and are protected against by the UPS. No action taken to resolve this, but we did look for breaker box or outlet box mis-wirings. Install new shutter battery -- We previously ordered and have now installed a new battery (car-battery size) for the dome shutter. It sits on a tray in the dome. In addition, cleaned up the cabling between the battery and the shutter controller. There was also a fault in the controller. The designer of the controller installed 'protection' for low battery voltage (shutter will not open or close if battery voltage is low). We bypassed this feature because we think it potentially does more harm than good; it may prevent proper closing of the shutter at the end of an observing run. Checked the primary mirror -- dirty. We're planning for a by-hand cleaning of the mirror on another visit (slow and tedious work). Telescope shroud -- Re secured telescope shroud.
To-Do items. -- Return camera for repair. Select, purchase, and install new video cards in obs and wx computer. Select, purchase, and install two new monitors in observatory. Re-cable video from computers to monitors. Clean primary mirror
Upgrade to fiber optic Internet pending. Will go from 6 Mb/sec to 200 Mb/sec.
4/17/2021 -- Problem with ICRON box (USB extender) being intermittent. Replaced power supply and several cables. Attached spade lugs to wiring at Digital Logger switch. APC UPS firmware update failed due to bad information from APC tech. New cable ordered. Update firmware on Digital-Logger switches. Analyze shutter battery -- needs replacement (on order). Installed new keyboard on computer.
3/8/2021 -- The telescope system has been out of service for a while. There was a problem with a bearing in the RA drivetrain. This problem has been fixed and the telescope is operational. There is always a maintenance list and an upgrade list of things to do.
12/20/2020 -- The Madrona Peak Observatory telescope has been in operation for most of 2020, without intervention from the 'maintenance crew'. In December , 2020 the telescope went out of service pending resolution of a problem with the drive system. The telescope will not move in RA. Work is in progress to identify and fix the problem.
1/11/2020 -- Installed primary imaging camera (now repaired) on the telescope. Cleaned the field flattener lens between the camera and the 'scope. Replaced wiring to camera from power supply to reduce voltage drop. Camera draws significant current when cooling, and the rewiring keeps the voltage to the camera above the minimum requirement (reduces resistive losses). Changed power source for camera from rack supply to the power supply provided with the camera.
11/23/19 -- The primary imaging camera has been returned after repairs (10 months). Installation is pending. The ICRON box (USB extender) that provides connectivity between the imaging camera and the computer (and other connections) has been replaced due to failure in the original unit. The new ICRON required a different power supply, which was installed. To avoid ground loops the 12 VDC power and the ICRON power grounds were connected together and to the telescope frame. The secondary imaging camera, which has been in place since February, 2019 has functioned intermittently. Cleaning the electrical connections seems to have helped. Yet another ICRON USB extender to transmit weather data from the console to the computer is pending installation. Updated the Acronis backup software to latest version.
9/8/19 -- After a bit of work the communications between the Obs and the Wx computers to their associated UPS ethernet cards is working. This allows for automated shutdown of the system in the event of a power outage. System is unavailable, though, due to no communications between the imaging camera and the Obs computer.
7/28/19 -- Resolved problem with system not returning to 'ready' state after a power outage. A setting in the Digital-Logger switches had to be changed to force a return to previous conditions. The imaging camera was not communicating with the imaging software. This was the result of a failure of a USB port connected to the computer. Moving the camera to a different USB port solved the problem. Installed ethernet cards in the APC UPS units. Configured the cards to be active on the ranch network. Successfully tested a new (home grown) dome driver to correct problems associated with previous dome driver. Resolved problem with connection of USB cable between UPS and computers.
2/22/19 -- Installed two new UPS units (and new management system), a second Digital-Loggers power switch, new weather sensors (temp, humidity, wind, etc.). Configured new Unifi network switch. Checked function of observatory. Ok, but Apogee Aspen camera had moisture on CCD; replaced with backup camera. Aspen camera returned to manufacturer for repair. Backup camera installed and tested, but new calibration images need to be taken. Surveillance camera previously installed. Checked computer backups to confirm that they are working. Davis weather console not working; new one ordered.
12/17/18 -- Some of the problems at the observatory are resolved. Internet communications to the observatory computer is now available. A trip is scheduled to install new remote power switches and a new surveillance camera. We expect this effort will return the system to full functionality.
10/29/18 -- Plan to assess the situation at the observatory in a couple of weeks followed by remedial work (if needed).
10/12/18 -- Observatory maintenance needed. System computer has malfunctioned along with communications to the observatory. A repair trip is being scheduled.
5/17/18 -- Cleaned cloud sensor. Checked clearance of mirror cover and camera slot. Resolved problem with guider camera (not being seen by computer). O/S update to weather computer resolved problem with restart.
1/6/18 -- Replaced hard drive in weather monitoring computer with 500 GB SSD. Resolved problems with slow booting of computer and failure to restart after issuing 'restart' command. Updated BIOS, reset boot on power connected. Resolved problem with 'Share' used to communicate weather information.
1/6/18 -- Tested main imaging camera USB connection to determine if USB and ethernet cables can be connected simultaneously. They cannot.
1/6/18 -- Cleaned lens of all-sky camera.
1/6/18 -- Resolved problem with UPS communication with computer. (New cable)
1/6/18 -- Identified problem with monitor in observatory not working.
1/6/18 -- Reconfigured power switches to isolate ethernet switch.
1/6/19 -- Tested emergency shutdown script (more work needed)
12/1/17 -- Pending work -- replace hard drive in weather monitor computer
3/11/17 -- Re-install original camera Apogee Aspen after repair. Recalibrate guide camera
1/28/17 -- Create wiring block diagrams for observatory control systems
1/28/17 -- Recalibrate autoguider
1/28/17 -- Update focuser 'home' point
1/28/17 -- Clean filter on observatory air conditioner
1/27/17 -- Resolve network communications problem to observatory
1/27/17 -- Recalibrate autoguider.
1/27/17 -- Install new batteries in UPS units. Enable UPS monitoring.
12/9/16 -- Remove Apogee Aspen camera from telescope. Send to ANDOR for repair of leak in 'window' above imager chip. Install backup camera -- Apogee Alta
11/19/16 -- Install backup disks.
11/19/16 -- Acquire external disks for backups
11/19/16 -- Fix frost on imager problem -- camera returned to vendor for repair
11/19/16 -- Resolve problem with reboot on power up of observatory computer
11/19/16 -- Install new power supply for observatory WiFi access point
11/19/16 -- Photos for documentation of existing system configuration
11/19/16 -- Learn about ACP scheduler -- executed successfully
11/18/16 -- Activate pointing model for system
11/13/16 -- Fix all sky camera upload problem
11/13/16 -- Upgrade SkyX to version 10.5
11/12/16 -- First run of telescope by current operator; image of Ring Nebula taken. Pointing good but did not include pointing model.